OpenStudy (anonymous):

Simplify: -27i^3 + 43i^7 - 8i^2 + 15i^3 A) 8 - 31i B) 8 + 31i C) 23 - 15i D) 23 + 15i

OpenStudy (mathstudent55):

First, combine like terms.

OpenStudy (mathstudent55):

Then look at this: \(i^0 = 1\) \(i^1 = i\) \(i^2 = -1\) \(i^3 = -i\) \(i^4 = 1\) etc.

OpenStudy (anonymous):

oh ok

OpenStudy (mathstudent55):

@mathacomplice Can you combine like terms and show what you get?

OpenStudy (mathstudent55):

The only like terms are shown in red: \(\large \color{red}{-27i^3} + 43i^7 - 8i^2 + \color{red}{15i^3}\)

OpenStudy (anonymous):

so it'll be -12i^3?

OpenStudy (mathstudent55):

Right. Now you have this: \(= -12i^3 + 43i^7 - 8i^2\)

OpenStudy (mathstudent55):

Now use my table above, and replace i^3 by -i, i^2 by -1 and can you figure out what i^7 is equal to? To figure it out, just continue the pattern.

OpenStudy (anonymous):

i, right?

OpenStudy (mathstudent55):

No. Notice that i^4 is the same as i^0. Once you reach i^4, use i^0 and continue from there.

OpenStudy (anonymous):

oh ok

OpenStudy (mathstudent55):